# arXiv:2603.01316v1[eess.AS] 01 Mar 2026

### v1

@@ Abstract
This paper investigates the use of relative cues for text-based target speech extraction (TSE). We first provide a theoretical justification for relative cues from the perspectives of human perception and label quantization, showing that relative cues preserve fine-grained distinctions often lost in absolute categorical representations. Building on this analysis, we propose a two-stage TSE framework, in which a speech separation model generates candidate sources, followed by a text-guided classifier that selects the target speaker based on embedding similarity. Using this framework, we train two separate classification models to evaluate the advantages of relative cues over independent cues in terms of both classification accuracy and TSE performance. Experimental results demonstrate that (i) relative cues achieve higher overall classification accuracy and improved TSE performance compared with independent cues, (ii) the two-stage framework substantially outperforms single-stage text-conditioned extraction methods on both signal-level and objective perceptual metrics, and (iii) certain relative cues (language, gender, loudness, distance, temporal order, speaking duration, random cue and all cue) can surpass the performance of an audio-based TSE system. Further analysis reveals notable differences in discriminative power across cue types, providing insights into the effectiveness of different relative cues for TSE.
@@ I Introduction
Subsequent developments incorporated pretrained speaker encoders, such as D-Vectors[55], X-Vectors[51], and ECAPA-TDNN[8], to obtain robust speaker embeddings, decoupling speaker modeling from the extraction network. Moreover, multi-modal approaches integrating audio, video cues have gained attention[37, 26, 54, 58], especially for audiovisual datasets like AVSpeech[11] or VoxCeleb[35]. These methods align visual speech movements or lip information with audio stream[3, 25], improving robustness in visually grounded TSE scenarios.
Despite these advancements, several challenges remain, including robustness to unseen speakers, limited availability of clean target references in real-world scenarios, and mismatch between the enrollment speech and the mixture input regarding recording conditions, speaking style, or acoustic environment. Beyond acoustic/visual-based guidance, recent research has explored text-guided extraction. These models use a text query (prompt) to describe speech attributes of target speaker as a condition to guide the extraction process of the desired speaker’s speech signal in the mixture, enabling extraction when no enrollment utterance of target speaker is available. Compared to other cues, text offers greater accessibility, enabling users to provide instructions without recording audio or video. It also provides greater flexibility, as it can convey a wide range of target speaker characteristics, including linguistic cues (e.g., speech transcription), paralinguistic cues (e.g., gender), and acoustic cues (e.g., loudness of speech). In addition, text offers stronger privacy by avoiding the transmission or storage of personal voice data.
Hao et al. were the first to successfully introduce text-based TSE that used a LLM for extracting semantic information from text queries and integrating it into a speech separation network [16]. Subsequently, Huo et al. proposed a flexible text-guided TSE framework and introduced the TextrolMix dataset, which provides rich textual descriptions of speech attributes, including speaker identity, emotion, pitch level, gender, accent, and speaking rate [17]. Similarly, Seki et al. developed a TSE model conditioned on text prompts describing paralinguistic and non-linguistic attributes such as gender, pitch, speaking speed, and loudness, demonstrating that it can overcome the limitations of existing audio-based TSE methods [46].
@@
To overcome this limitation, our prior work introduced inter-speaker relative cues, which explicitly describe differences between the target and interfering speakers [7]. By formulating attributes relatively rather than as speaker-specific categories, these cues preserve fine-grained distinctions that are critical for target speaker extraction. To study these cues, we constructed a two-speaker mixture dataset spanning a wide range of attributes, including pitch level, pitch range, speaking rate, speaking duration, temporal order, age, loudness, speaker-to-microphone distance, language, gender, transcription and emotion [7].
This approach is motivated by human perception: people perceive continuous-valued speech attributes primarily through comparative listening, rather than by measuring precise physical quantities. Instead of relying on exact values—such as fundamental frequency ($F_{0}$) in Hertz or speaking rate in syllables per minute—listeners judge whether one speaker sounds higher or lower in pitch, or faster or slower in speaking rate than another [53, 33, 32]. Conventional TSE systems that rely on absolute, discretized labels fail to capture these relative distinctions. Subtle but perceptually meaningful differences between speakers can be lost when multiple speakers fall into the same category (e.g., both labeled as “fast” speaking rate). By explicitly encoding inter-speaker comparisons—such as “Speaker A has a faster speaking rate than Speaker B”—relative attribute cues preserve these contrasts, providing more discriminative and perceptually aligned conditioning for target speaker extraction.
@@
  • Theoretical Justification of Relative Cues: We provide a theoretical analysis of relative cues, which accounts for human perception and label quantization, offering a principled explanation of their advantages over speaker independent cues, which we hereafter refer to as independent cues throughout the remainder of this paper.
  • Proposed Two-Stage Text-Guided TSE Method: In our previous text-based TSE system—similar to conventional audio-based approaches—the conditioning-and-extraction paradigm is prone to speaker confusion [62, 28, 34, 10] during inference, leading to degraded performance. Motivated by enrollment-audio-based two-stage TSE approach [10], and to improve upon our previous single-stage system, we adopt a two-stage framework in this work: first, performing speech separation, and then using a classification model to select the target speaker based on inter-speaker relative cues embedded in text prompts. This approach not only mitigates the speaker confusion inherent in single-stage TSE systems but also provides interpretability, allowing insight into the relative effectiveness of different cues through analysis of the classification accuracy for each cue.
  • Empirical Evaluation of the Advantages of Relative Cues: Based on the proposed two-stage framework, we evaluate the impact of using relative versus independent cues on overall target speech classification and the corresponding TSE performance, focusing on continuous-valued attributes. This provides an empirical perspective on the advantages of relative cues over independent cues.
@@ II Inter-Speaker Relative Cues
    appearance time, defined as the onset time of a speaker’s speech within the mixture.
@@
    transcription content, corresponding to the linguistic content of the utterance in textual form.
We derive each inter-speaker relative cues by comparing a single speech attribute between the target speaker and the interfering speaker in a two-speaker mixture. Each relative cue encodes the difference along one specific attribute only. Independent cues, in contrast to relative cues, are derived solely from the attributes of an individual speaker—either the target or the interfering speaker—without accounting for inter-speaker relationships.
@@ II-A Independent Cues
where $\mathcal{Q}_{i}(\cdot)$ denotes the quantization operator for the $i$-th attribute, $k_{i}$ is the number of discrete categories for that attribute, and $\mathit{A}_{i,1},\dots,\mathit{A}_{i,k_{i}}$ represent the discrete categories (e.g., low, normal, and high for mean $F_{0}$).
The corresponding independent cue category for attribute $i$ is given by $\tilde{x}_{i}$. This mapping from continuous-valued attribute values $x_{i}$ to discrete independent cue categories $\tilde{x}_{i}$ is illustrated in Fig. 1(a). Due to the coarse nature of the categories, perceptually noticeable differences in the original continuous-valued attribute may be mapped to the same discrete category (as illustrated by the shaded region in Fig. 1(a)), in which case the attributes of two speakers cannot be distinguished from each other.
(a) Independent Cue
(b) Relative Cue for Absolute Difference
(c) Relative Cue for Percentage Difference
Figure 1: Illustration of quantization schemes for continuous-valued attributes. (a) Independent cues: the attribute value space of the target and interfering speakers, denoted by $x_{i}^{(\mathit{tar})}$ (y-axis) and $x_{i}^{(\mathit{inf})}$ (x-axis), is partitioned into a small number of coarse categories (three shown as in the example). Shaded regions indicate attribute values of the two speakers that are quantized into the same category, in which case the two speakers cannot be distinguished based on the attribute. (b) Relative cues for absolute differences: attribute values of the target and interfering speakers are compared with absolute difference. The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold. Regions outside the shaded area represent perceptually distinct absolute differences. (c) Relative cues for percentage differences: attribute values of the target and interfering speakers are compared with percentage difference. The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold. Because the comparison is based on relative differences, the Similar region gradually expands from the zero point. Regions outside the shaded area represent perceptually distinct relative differences.
@@ II-B Relative Cues
In contrast, relative cues categorize continuous-valued attributes into groups (e.g., slower, similar, and faster for speaking rate) by comparing the attribute values of the two speakers. For the attributes RMS energy (dB), speaker-to-microphone distance, speaker age, and appearance time in the mixture, the difference is expressed as the absolute difference
$\Delta x_{i}^{|.|}=x_{i}^{(\mathrm{tar})}-x_{i}^{(\mathrm{inf})}.$ (2)
These attributes correspond to interpretable physical or temporal quantities whose perceptual relevance is largely determined by their absolute values. Changes in these quantities can be directly perceived and interpreted by human listeners without reference to a baseline. For attributes such as mean $F_{0}$, $F_{0}$ span, speaking rate, and speaking duration, we employ the percentage difference
@@
For notational convenience, we introduce the variable $\Delta x_{i}^{(*)}\in\{\Delta x_{i}^{|.|},\Delta x_{i}^{\%}\}$ to denote the attribute-specific difference measure. This measure is then compared against the predefined threshold $\theta_{i}$ of each attribute, which are shown in Table I.
TABLE I: Continuous-Valued Speech Attributes with Our Defined Thresholds for Relative Cues and Motivated Speech Perceptual Studies
Speech Attribute Threshold ($\theta_{i}$) Motivated Findings
RMS energy (dB) 3 dB Perception of relative loudness differences in speech is mediated by changes in loudness-related acoustic features, such as excitation strength derived from the glottal source [47]. McShefferty et al. [30] report that the average just-noticeable difference (JND) in speech-to-noise ratio is approximately 3 dB. Signal-to-interference ratio (SIR), computed based on RMS energy over active speech regions, can serve as an approximation of the perceived loudness difference between target and interfering speech.
@@
where $\mathit{R}_{i,1}$, $\mathit{Similar}$, and $\mathit{R}_{i,2}$ denote relational categories corresponding, for example, to longer, similar, and shorter speaking duration relative to the interfering speaker. This mapping is illustrated in Fig. 1(b) and Fig. 1(c). The distribution of these relative cues under different threshold settings in the test set (based on the simulated dataset described in Section IV) are presented in Table II. The total number of samples is almost balanced across cue types, except for speaker age, for which labeled samples are limited in our dataset.
TABLE II: Sample counts of relative categories for each continuous attribute in the test set, based on the defined thresholds.
Continue-valued Attribute R1 Similar R2
@@ II-D Text Prompt Generation
Inter-speaker relative cues are formulated as natural-language descriptions (text prompts) that explicitly characterize the relative differences between the target and interfering speakers. These prompts are used for both training and evaluation of the TSE systems. Each text prompt is generated by populating predefined templates with one or more relative cue descriptors, following the same approach as [7]. The templates incorporate action-oriented verbs such as extract, separate, or isolate, combined with descriptive phrases that emphasize the distinguishing characteristics of the target speaker. For example, a cue vector indicating (female, higher pitch, faster speaking rate) can be verbalized as: “Please extract the female speaker with a higher pitch and faster speaking rate.” We also employed the GPT-4o-mini model to automatically rephrase the templates into more generic prompts. Although these diversified prompts increased linguistic variability, they resulted in lower performance. Therefore, we adopt the original template-based prompts in our experiments. Since relative cues are mixture-specific, the target speaker in a given mixture may be specified by one or multiple relative cues.
Table IV reports the number of mixtures in each subset that contain a given relative cue. Since a single mixture may include multiple cues, the counts for individual cues do not sum to the total number of mixtures. As shown in the table, the target speaker in each mixture can be specified using different types of prompts. The Random Cue type consists of a combination of randomly selected relative cues, where the number of selected cues ranges from 2 to (n1)(n-1) and $n$ denotes the total number of relative cues in the complete set. In contrast, the All Cue type includes all $n$ relative cues for a given mixture. This procedure yields a diverse set of natural-language descriptions that accurately reflect the underlying attribute differences between speakers.

### v2

# Inter-Speaker Relative Cues for Two-Stage Text-Guided Target Speech Extraction

@@ Abstract
This paper investigates the use of relative cues for text-based target speech extraction (TSE). We first provide a theoretical justification for relative cues from the perspectives of human perception and label quantization, showing that relative cues preserve fine-grained distinctions that are often lost in absolute categorical representations for continuous-valued attributes. Building on this analysis, we propose a two-stage TSE framework in which a speech separation model first generates candidate sources, followed by a text-guided classifier that selects the target speaker based on embedding similarity. Within this framework, we train two separate classification models to evaluate the advantages of relative cues over independent cues in case of continuous-valued attributes, considering both classification accuracy and TSE performance. Experimental results demonstrate that (i) relative cues achieve higher overall classification accuracy and improved TSE performance compared with independent cues; (ii) the proposed two-stage framework substantially outperforms single-stage text-conditioned extraction methods on both signal-level and objective perceptual metrics; and (iii) several relative cues, including language, loudness, distance, temporal order, speaking duration, random cues, and all cues, can even surpass the performance of an enrollment-audio-based TSE system. Further analysis reveals notable differences in discriminative power across cue types, providing insights into the effectiveness of different relative cues for TSE.
@@ I Introduction
Subsequent developments incorporated pretrained speaker encoders, such as D-Vectors[49], X-Vectors[46], and ECAPA-TDNN[8], to obtain robust speaker embeddings, decoupling speaker modeling from the extraction network. Moreover, multi-modal approaches integrating audio and video cues have gained attention[35, 24, 48, 52], especially for audiovisual datasets like AVSpeech[11] or VoxCeleb[34]. These methods align visual speech movements or lip information with audio stream[3, 23], improving robustness in visually grounded TSE scenarios.
Despite these advancements, several challenges remain, including robustness to unseen speakers, limited availability of clean target references in real-world scenarios, and mismatch between the enrollment speech and the mixture input regarding recording conditions, speaking style, or acoustic environment. Beyond acoustic/visual-based guidance, recent research has explored text-guided extraction. These models use a text query (prompt) describing the speech attributes of a target speaker as a condition to guide the extraction of the desired speaker’s speech signal from a mixture, enabling extraction when no enrollment utterance of the target speaker is available. Compared to other cues, text offers greater accessibility, enabling users to provide instructions without recording audio or video. It also provides greater flexibility, as it can convey a wide range of target speaker characteristics, including linguistic cues (e.g., speech transcription), paralinguistic cues (e.g., gender), and acoustic cues (e.g., loudness of speech). In addition, text offers stronger privacy by avoiding the transmission or storage of voice data or other media.
Hao et al. were the first to successfully introduce text-based TSE that used a LLM for extracting semantic information from text queries and integrating it into a speech separation network [15]. Subsequently, Huo et al. proposed a flexible text-guided TSE framework and introduced the TextrolMix dataset, which provides rich textual descriptions of speech attributes, including speaker identity, emotion, pitch, gender, accent, and speaking rate [16]. Similarly, Seki et al. developed a TSE model conditioned on text prompts describing paralinguistic and non-linguistic attributes such as gender, pitch, speaking speed, and loudness, demonstrating that it can overcome the limitations of existing audio-based TSE methods [42].
@@
To overcome this limitation, our prior work introduced inter-speaker relative cues, which explicitly describe differences between the target and interfering speakers [7]. By formulating attributes relatively rather than as speaker-specific categories, these cues preserve fine-grained distinctions that are critical for target speaker extraction. To study these cues, we constructed a two-speaker mixture dataset spanning a wide range of attributes, including pitch, pitch range, speaking rate, speaking duration, temporal order, age, loudness, speaker-to-microphone distance, language, gender, transcription and emotion [7].
This approach is motivated by human perception: people perceive continuous-valued speech attributes primarily through comparative listening, rather than by measuring precise physical quantities. Instead of relying on exact values, such as fundamental frequency ($F_{0}$) in Hertz or speaking rate in syllables per minute, listeners judge whether one speaker sounds higher or lower in pitch, or faster or slower in speaking rate than another [31, 30]. Conventional TSE systems that rely on absolute, discretized labels fail to capture these relative distinctions. Subtle but perceptually meaningful differences between speakers can be lost when multiple speakers fall into the same category (e.g., both labeled as “fast” speaking rate). By explicitly encoding inter-speaker comparisons, such as “Speaker A has a faster speaking rate than Speaker B”, relative attribute cues preserve these contrasts, resulting in more discriminative and perceptually aligned conditioning for target speaker extraction. Although this study restricts mixtures to two speakers, the relative cue formulation can be naturally extended to mixtures with more speakers. For instance, in mixtures with three or more speakers, cues such as “fastest speaking rate” or “second-fastest speaking rate” can be used to specify the target speaker. For discrete-valued speech attributes, listeners identify the target speaker by comparing the attribute category with that of the interfering speaker, further illustrating the principle of relative cues rather than speaker-independent cues.
@@
  • Theoretical Justification of Relative Cues: We provide a theoretical analysis of relative cues, which accounts for human perception and label quantization, offering a principled explanation of their advantages over speaker-independent cues, which we hereafter refer to as independent cues throughout the remainder of this paper.
  • Proposed Two-Stage Text-Guided TSE Method: Similar to conventional audio-based approaches, our previous text-based TSE system follows a conditioning-and-extraction paradigm that is prone to speaker confusion during inference [55, 26, 33, 10], leading to degraded performance. Motivated by enrollment-audio-based two-stage TSE approach [10], and to improve upon our previous single-stage system, we adopt a two-stage framework in this work: first, performing speaker-independent speech separation, and then using a classification model to select the target speaker based on inter-speaker relative cues embedded in text prompts. This approach not only mitigates the speaker confusion inherent in single-stage TSE systems but also provides interpretability, allowing insight into the relative effectiveness of different cues through analysis of the classification accuracy for each cue.
  • Empirical Evaluation of the Advantages of Relative Cues: Based on the proposed two-stage framework, we evaluate the impact of relative versus independent cues on overall target speaker classification and the corresponding TSE performance for comparable continuous-valued attributes. This analysis provides an empirical perspective on the advantages of relative cues over independent cues.
@@ II Inter-Speaker Relative Cues
(a) Independent Cue
(b) Relative Cue for Direct Difference
(c) Relative Cue for Percentage Difference
Figure 1: Illustration of quantization schemes for continuous-valued attributes. (a) Independent cues: the attribute value space of the target and interfering speakers, denoted by $x_{i}^{(\mathit{tar})}$ (y-axis) and $x_{i}^{(\mathit{inf})}$ (x-axis), is partitioned into a small number of coarse categories (three shown as in the example). Shaded regions indicate attribute values of the two speakers that are quantized into the same category, in which case the two speakers cannot be distinguished based on the attribute. (b) Relative cues for direct differences: attribute values of the target and interfering speakers are compared with direct difference. The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold. Regions outside the shaded area represent perceptually distinct differences. (c) Relative cues for percentage differences: attribute values of the target and interfering speakers are compared with percentage difference. The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold. Because the comparison is based on relative differences, the Similar region gradually expands from the zero point. Regions outside the shaded area represent perceptually distinct relative differences.
@@
    appearance time, defined as the onset time of the first occurrence of a speaker’s speech in the mixture.
@@
    transcription, corresponding to the linguistic content of the utterance in textual form.
We derive each inter-speaker relative cue by comparing a single speech attribute between the target speaker and the interfering speaker in a two-speaker mixture. Each relative cue encodes the difference along one specific attribute only. Independent cues, in contrast to relative cues, are derived solely from the attributes of an individual speaker—either the target or the interfering speaker—without accounting for inter-speaker relationships.
@@ II-A Independent Cues
where $\mathcal{Q}_{i}(\cdot)$ denotes the quantization operator for the $i$-th attribute, $k_{i}$ is the total number of discrete categories for that attribute, and $\mathit{A}_{i,1},\dots,\mathit{A}_{i,k_{i}}$ denote the corresponding categories (e.g., low, normal, and high for mean $F_{0}$).
The independent cue category for attribute $i$ is given by $\tilde{x}_{i}$. The mapping from continuous-valued attribute values $x_{i}$ to discrete independent cue categories $\tilde{x}_{i}$ is illustrated in Fig. 1(a). Due to the coarse nature of the categories, perceptually noticeable differences in the original continuous-valued attribute may be mapped to the same discrete category (as illustrated by the shaded region in Fig. 1(a)), in which case the attributes of two speakers cannot be distinguished from each other.
@@ II-B Relative Cues
In contrast, relative cues categorize continuous-valued attributes into groups (e.g., slower, similar, and faster for speaking rate) by comparing the attribute values of the two speakers. For the attributes RMS energy (dB), speaker-to-microphone distance, speaker age, and appearance time in the mixture, the difference is expressed as the direct difference
$\Delta x_{i}^{(.)}=x_{i}^{(\mathrm{tar})}-x_{i}^{(\mathrm{inf})}.$ (2)
These attributes correspond to interpretable physical or temporal quantities whose perceptual relevance is largely determined by differences in their values. In human perception, changes in stimulus magnitude can be detected through difference thresholds, allowing listeners to perceive relative differences in such attributes between competing sources. Consequently, variations in these quantities can be directly perceived and interpreted without requiring an explicit baseline reference [32, 6]. For attributes such as mean $F_{0}$, $F_{0}$ span, speaking rate, and speaking duration, we employ the percentage difference
@@
For notational convenience, we introduce the variable $\Delta x_{i}^{(*)}\in\{\Delta x_{i}^{(.)},\Delta x_{i}^{\%}\}$ to denote the attribute-specific difference measure. This measure is then compared against the predefined threshold $\theta_{i}$ of each attribute, which are shown in Table I.
TABLE I: Continuous-Valued Speech Attributes with Our Defined Thresholds for Relative Cues and Motivating Speech Perceptual Studies
Speech Attribute Threshold ($\theta_{i}$) Findings from Motivating Speech Perception Literature
RMS energy (dB) 3 dB Perception of relative loudness differences in speech is mediated by changes in loudness-related acoustic features, such as excitation strength derived from the glottal source [43]. McShefferty et al. [28] report that the average just-noticeable difference (JND) in speech-to-noise ratio is approximately 3 dB.
@@
where $\mathit{R}_{i,1}$, $\mathit{Similar}$, and $\mathit{R}_{i,2}$ denote relational categories corresponding, for example, to longer, similar, and shorter speaking duration relative to the interfering speaker. This mapping is illustrated in Fig. 1(b) and Fig. 1(c). The distribution of these relative cues under different threshold settings in the test set (based on the simulated dataset described in Section IV) is presented in Table II. The total number of samples is nearly balanced across cue types, except for speaker age, for which labeled samples are limited in our dataset.
TABLE II: Sample counts of relative categories for each continuous attribute in the test set based on defined thresholds
Continuous-valued Attribute R1 Similar R2
@@ II-D Text Prompt Generation
Inter-speaker relative cues are formulated as natural language descriptions (i.e., text prompts) that explicitly capture the differences between the target and interfering speakers, and are used for both training and evaluation of the TSE systems. For each mixture, a target speaker is specified based on the dataset annotations (see Section IV), and a corresponding prompt is generated by instantiating a predefined template with one or more relative cues. Specifically, one prompt is generated for each available individual cue, as well as for combinations of cues, either randomly selected or comprising all available cues (see details below). In the first experimental setting (Section IV-C), cue categories are chosen to ensure clear inter-speaker contrast, thereby providing sufficient discriminability for target speaker extraction. In contrast, the second experimental setting (Section IV-D) adopts a different definition of cue categories, allowing labels such as “same” or “similar” during testing. The template structure used for prompt generation follows the implementation described in the code [^2] https://github.com/daiwangsnr/Inter-Speaker-Relative-Cues-for-Text-Guided-Target-Speech-Extraction/blob/main/prompt_template.py, incorporating action-oriented verbs such as extract, separate, or isolate, along with descriptive phrases that emphasize the distinguishing characteristics of the target speaker. For example, a cue vector such as (female, higher pitch, faster speaking rate) can be verbalized as: “Please extract the female speaker with a higher pitch and faster speaking rate.”
Prompts are generated in three configurations for each mixture:
  • Individual cues: A separate prompt is generated for each available cue.
  • Random cues: If at least three cues are available, a subset is randomly sampled for each mixture, with size $2$ to n1n-1 (where $n>3$ is the total number of available cues), and used to generate a prompt. This configuration enables robust evaluation under partial and randomly combined cues.
  • All cues: If at least two cues are available, a prompt is generated that incorporates all cues, enabling evaluation under fully combined cue conditions.
We also experimented with GPT-4o-mini to automatically rephrase templates for greater linguistic diversity. Although this increased variation, performance slightly decreased; thus, template-based prompts are retained in our experiments. Table III (Section IV) reports the number of mixtures associated with each relative cue in each subset. Some cue names differ from their corresponding speech attributes (e.g., loudness [^3] In this work, “loudness” refers to RMS energy and does not strictly correspond to its conventional perceptual definition vs. RMS energy, age vs. speaker age, pitch vs. mean $F_{0}$, pitch range vs. $F_{0}$ span, distance vs. speaker-to-microphone distance, temporal order vs. appearance time, and emotion vs. emotional state). These differences arise because some cue names are abbreviated, while others are expressed in more perceptually meaningful terms. Since not all mixtures include every cue (i.e., the target speaker is not necessarily characterized by all possible cues), the counts for individual cues do not sum to the total number of mixtures.